A solid conducting sphere of radius 2.00cm has a charge of 8.00 uC. A conducting spherical shell of inner radius 4.00 cm and outer radius 5.00 cm is concentric with the solid sphere and has a charge of -4.00 uC. Find the electric field at (a) r=1.00 cm (b) r=3.00 cm (c)r =4.50 cm and (d)r=7.00 cm from the centre of this charge configuration.
1
Expert's answer
2020-03-25T11:16:26-0400

a)


"E=0\frac{N}{C}"

b)


"E=\frac{kq_1}{r^2}=\frac{8.99\cdot 10^9\cdot 8\cdot 10^{-6}}{0.03^2}=7.99\cdot 10^7\frac{N}{C}"

c)


"E=0\frac{N}{C}"

d)


"E=k\frac{q_1}{r^2}=8.99\cdot 10^9\frac{(8-4)\cdot 10^{-6}}{0.07^2}=7.34\cdot 10^6\frac{N}{C}"